# Magnetically-Driven Neutron-Rich Ejecta Unleashed: Global 3D Neutrino-GRMHD Simulations of Collapsars Reveal the Conditions for $r$-process Nucleosynthesis

### v1

@@ Abstract
Collapsars – rapidly rotating stellar cores that form black holes (BHs) – can power gamma-ray bursts (GRBs) and are proposed to be key contributors to the production of heavy elements in the Universe via the rapid neutron capture process ($r$-process). Previous neutrino-transport collapsar simulations have been unable to unbind neutron-rich material from the disk. However, these simulations have not included magnetic fields or the BH, both of which are essential for launching mass outflows. We present $\nu$h-amr, a novel neutrino-transport general relativistic magnetohydrodynamic ($\nu$GRMHD) code, which we use to perform the first 3D $\nu$GRMHD collapsar simulations. We find a self-consistent formation of a disk with initially weak magnetic flux, resulting in a low accretion speed and leaving sufficient time for the disk to neutronize. However, once substantial magnetic flux accumulates near the BH, it becomes dynamically important, leading to a magnetically arrested disk that unbinds some of the neutron-rich material. The strong flux also accelerates the accretion speed, preventing further disk neutronization. The neutron-rich disk ejecta collides with the infalling stellar gas, generating a shocked cocoon with an electron fraction, $Y_{\text{e}}\gtrsim 0.2$. Continuous mixing between the cocoon and neutron-poor stellar gas incrementally raises the outflow $Y_{\text{e}}$, but the final $r$-process yield is determined earlier at the point of neutron capture freeze-out. Our models require extreme magnetic fluxes and mass accretion rates to eject neutron-rich material ($Y_{\text{e}}\lesssim 0.3$), implying very high $r$-process ejecta masses $M_{\text{ej}}\lesssim{}M_{\odot}$. Future work will explore under what conditions more typical collapsar engines become $r$-process factories.
@@ 1 Introduction
Generally, the $r$-process nucleosynthesis channel requires neutron-rich outflows. Accretion disks that form during the collapse of rapidly rotating massive stars exhibit similar density and temperature conditions as those found in post-merger disks. Moreover, the larger size and lifetime of collapsar disks render them more massive than post-merger disks, suggesting their potential as the major source of the Universe’s heavy $r$-process elements (Siegel et al., 2019). However, the stellar infall, which forms collapsar disks, likely obstructs the ejection of neutron-rich disk outflows. Describing this complex, non-linear behavior necessitates numerical simulations to accurately model the formation of the collapsar disk, the synthesis of heavy elements, and the propagation of outflows through the collapsing star.
@@
The collapse of a massive stellar core initially forms a proto-NS. The $r$-process nucleosynthesis can happen in NS-outflows, which are a combination of neutrino-driven (e.g., Thompson et al., 2004; Metzger et al., 2007) and magnetically-driven (Kotake et al., 2004; Ardeljan et al., 2005; Moiseenko et al., 2006; Obergaulinger et al., 2006; Obergaulinger & Aloy, 2020, 2021; Mösta et al., 2014; Kuroda et al., 2020; Aloy & Obergaulinger, 2021). Neutrino-heated winds can facilitate favorable conditions for the $r$-process only for rapidly rotating, strongly magnetized proto-NSs, thanks to the shorter expansion timescales (Thompson et al., 2004; Metzger et al., 2007, 2008b; Vlasov et al., 2014, 2017; Desai et al., 2022; Prasanna et al., 2023). Studies of magnetorotational SNe (MR-SNe) show that they require rapid rotation and strong magnetic fields to launch powerful magnetically-driven outflows, or jets (Nishimura et al., 2006; Winteler et al., 2012; Nishimura et al., 2015; Tsujimoto & Nishimura, 2015; Halevi & Mösta, 2018; Mösta et al., 2018; Reichert et al., 2021, 2023). Strong outflows help to unbind neutron-rich matter and drive it at high velocities to avoid neutrino irradiation from the proto-NS, leading to robust $r$-process nucleosynthesis. Weaker initial magnetic field strengths lead to weaker outflows or delayed jet launching, and therefore result in the $r$-process element production only up to the 2nd peak (Nishimura et al., 2006; Winteler et al., 2012; Nishimura et al., 2015). Moreover, effects like current-driven kink instabilities and misalignment between the magnetic and rotation axes can also reduce the ability of MR-SNe to contribute to heavy $r$-process enrichment (Halevi & Mösta, 2018; Mösta et al., 2018).
@@
Alternatively, neutron-rich accretion disks can also form around BHs, i.e., in collapsars (Woosley, 1993; MacFadyen & Woosley, 1999; MacFadyen et al., 2001; Fujimoto et al., 2004; Pruet et al., 2004; Kohri et al., 2005; Surman et al., 2006; Woosley & Bloom, 2006). Growing computational power and advances in computational methods have enabled the numerical simulations of the $r$-process nucleosynthesis in this context. Earlier calculations showed neutron-rich material production and successful $r$-process operation in collapsar jets (Fujimoto et al., 2007; Ono et al., 2012; Nakamura et al., 2013). More recently, simulations modeled the magnetized evolution of an idealized isolated preset torus in hydrostatic equilibrium, i.e., without including the progenitor star (Siegel et al., 2019; Miller et al., 2020). These works show that collapsar disks may become neutron-rich. However, it remains unclear whether the behavior of these idealized tori resembles that of disks self-consistently formed during the stellar collapse and whether the neutron-rich elements can overcome the ram pressure of the infalling stellar gas and escape out of the collapsing star. More recently, simulations included the collapsing star for the first time, which hindered the ejection of the heavy $r$-process elements (Just et al., 2022; Fujibayashi et al., 2023; Dean & Fernández, 2024a, b). However, these simulations lacked important physics: they reduced the dimensionality to 2D and excluded magnetic effects. The absence of magnetic fields does not allow magnetized outflows to form via, e.g., the Blandford-Znajek (BZ; Blandford & Znajek, 1977) and Blandford-Payne–like (Blandford & Payne, 1982) processes, which may potentially carry away some of the neutron-rich elements.
@@
In this Letter, we present the first 3D neutrino-transport general-relativistic magnetohydrodynamic ($\nu$GRMHD) collapsar simulations. Our simulations use a two-moment (M1) neutrino transport scheme. We find that including both the self-consistent disk formation in the collapsing stellar gas and subsequent magnetic launching of collimated disk outflows is crucial for the development of neutron-rich ejecta in collapsars. We begin by discussing the theoretical requirements for forming and ejecting neutron-rich material in §2. Using these arguments, we motivate our numerical setup in §3. We present our simulation results in §4 and discuss their implications in §5.
@@ 2 How To Make $r$-process Ejecta
Here, we outline the conditions under which collapsar disks form, generate neutron-rich material, and subsequently eject it. An accretion disk forms if the collapsing stellar gas possesses specific angular momentum, $l(r)$, exceeding that at the innermost stable circular orbit (ISCO), $r_{\rm ISCO}$:
@@
where $r$ is the spherical polar radius, $\,{M_{\rm BH}}$ is the BH mass, and for simplicity, we use a non-relativistic approximation for the specific angular momentum at the ISCO. There are theoretical indications that collapsar progenitors feature rapid rotation, which results in early disk formation (e.g., Gottlieb et al., 2024).
A necessary, but not sufficient, condition for the $r$-process is that the disk must become neutron-rich. The degree of neutron-richness is inversely correlated with the electron fraction $\,{Y_{\rm e}}=n_{\rm p}/(n_{\rm p}+n_{\rm n})$, where $n_{\rm p}$ and $n_{\rm n}$ are the proton and neutron number densities respectively. For weak interactions to drive the disk material neutron-rich ($\,{Y_{\rm e}}<0.5$), it must become dense and mildly degenerate through neutrino cooling (e.g., Beloborodov 2003). Specifically, when the viscous heating rate is counterbalanced by neutrino cooling, the disk transitions from a thick advective state to an efficiently neutrino-cooled regime (e.g., Di Matteo et al. 2002), in which it can become notably dense and neutron-rich. Satisfying this “ignition” condition requires a minimum mass accretion rate at the ISCO (Chen & Beloborodov, 2007; Metzger et al., 2008a, b),
@@
where $\alpha_{\rm eff}$ is the effective viscosity, and $K_{\rm ign}$ is a numerical prefactor that depends on the BH spin $a_{\rm BH}$: $K_{\rm ign}=0.07$ for $a_{\rm BH}=0$ and $K_{\rm ign}=0.02$ for $a_{\rm BH}=0.95$.
@@
and consequently becomes neutron-rich. At the ignition radius, $R=R_{\rm ign}$, the neutrino cooling and viscous heating balance each other. In classical accretion disk theory (e.g., for a standard Shakura & Sunyaev 1973 disk), a single viscosity coefficient, $\alpha_{\rm eff}$, governs both the viscous heating and accretion. However, in the presence of strong magnetic fields, large-scale magnetic torques can render the accretion more efficient and decouple it from the heating (Blandford & Payne, 1982; Ferreira & Pelletier, 1995). In practice, for strongly magnetized disks, the accretion timescale is a factor of a few shorter than the viscous heating timescale (Manikantan et al., 2024). Therefore, we can define the ignition radius as the radius where the neutrino cooling timescale, $t_{\rm cool}$, matches the accretion timescale, $t_{\rm acc}$:
@@
Here, the cooling timescale is (Metzger et al., 2008b):
@@
where $\dot{N}_{e^{-}p}$ is the rate of electron capture on protons and $\langle\cdot\rangle^{\theta,\varphi}_{\rho}$ is a density-weighted average over angles $\theta,\varphi$, which for quantity $X$ is:
@@
Whereas disk winds may eject neutron-rich material in compact binary mergers (e.g., Fernández & Metzger, 2013; Siegel & Metzger, 2017; Radice et al., 2018), they might be too weak to unbind the neutron-rich material in the inner parts of collapsar disks (e.g., Dean & Fernández, 2024a). However, powerful relativistic BZ jets, associated with GRBs, can explode the star and eject the neutron-rich material out of the collapsing star. The formation of such jets becomes possible once enough poloidal (i.e., pointing in the $r$- and $\theta$-directions) magnetic flux accumulates on the BH to overcome the ram pressure of the infalling gas. Namely, when the Alfvén velocity becomes comparable to the free-fall velocity, the jet can emerge from the BH event horizon (Komissarov & Barkov, 2009; Gottlieb et al., 2023).
The jet launching criterion is determined by the dimensionless magnetic flux on the BH, defined as:
@@
where $\,{r_{\rm g}}\equiv GM_{\rm BH}/c^{2}$ is the BH gravitational radius. To infer the minimum value of $\phi$ for jet launching, we use estimates by Komissarov & Barkov (2009), where they found $\phi_{\mathrm{crit}}\sim 20$ in idealized simulations of jet launching for radial magnetic geometry (this corresponds to $\kappa_{\rm c}\sim 1.5$ in their work). In addition to powerful BH-powered outflows, magnetically arrested disks (MAD; Narayan et al., 2003; Tchekhovskoy et al., 2011) can also launch massive disk winds whose properties we study below. The accretion flow reaches the maximum magnetization when it goes MAD, which occurs at $\phi\approx 50$ (Tchekhovskoy et al., 2011).
We conclude that collapsars can eject $r$-process elements if: (1) they form an accretion disk with high mass accretion rates and efficient neutrino cooling, and (2) once the $r$-process elements are synthesized in the disk, dynamically important magnetic fields around the BH launch magnetized jets and winds that manage to unbind the neutron-rich material.
@@ 3.1 Initial conditions
Table 1: Collapsar model parameters. Leftmost column: model name, which follows [cp]B[0ws][A#] convention, where c or p indicates stellar density profile (core or power-law, respectively), 0, w, or s indicates the magnetic field strength (zero, weak, and strong field), and A# indicates BH spin (omitted for the fiducial value, $a_{\rm BH}=0.8$). The rest of the columns, from left to right: initial density slope, $\alpha_{\rm p}$; grid resolution, $N_{r}\times N_{\theta}\times N_{\phi}$; simulation duration, $t_{\rm f}$; maximum initial core magnetic field strength, $B_{\rm core}$; MAD onset time, $t_{\rm MAD}$; jet (outflow) launch time, $t_{\rm jet}$; $\dot{M}$ at these two times; jet power at the larger of these two times, $L_{\rm j}$; total mass of neutron-rich ($\,{Y_{\rm e}}<0.25$) material at $t_{\rm MAD}$, $M_{\rm nr}$. Parameter values common to all models: BH mass, $M_{\rm BH}=4\,{M_{\odot}}$; stellar envelope total mass $M_{\rm star}=70\,{M_{\odot}}$; stellar radius, $R_{\rm star}=4\times 10^{10}$ cm; magnetic core radius, $R_{\rm core}=10^{8}$ cm (such that $\mu=B_{\rm core}R_{\rm core}^{2}/2$ in Eq. 12).
@@
To determine the electron fraction evolution, we implement a novel two-moment (M1) neutrino transport scheme alongside the Helmholtz equation of state (EOS; see §3.4) into the GPU-accelerated code h-amr (Liska et al., 2022). Using the resulting $\nu$h-amr code, we conduct a suite of 3D neutrino-transport general relativistic magnetohydrodynamic ($\nu$GRMHD) simulations of collapsing stars, as outlined in Table 1. In the following, we adopt the units $G=c=1$, and use Heaviside-Lorentz units for magnetic fields and thereby absorbing the factor of $1/\sqrt{4\pi}$ into the definition of magnetic field strength, $B$.
We set up our collapsar simulations following Gottlieb et al. (2022a, b). We set the newly formed BH mass to $\,{M_{\rm BH}}=4\,\,{M_{\odot}}$ and its dimensionless spin magnitude to $\,{a_{\rm BH}}=0.8$. Our simulation incorporates a static metric such that the BH mass and spin do not evolve over the course of the simulation. The BH is embedded in a stripped-envelope (e.g., Wolf-Rayet-like) star of mass [^1] As we will show, at the relevant simulation times (MAD onset timescale), the accreted mass is $\lesssim 10\,\,{M_{\odot}}$. This implies that a slightly more extended core with a steeper density profile outside will produce the same accretion rates with $M_{\star}\approx 20\,\,{M_{\odot}}$. $M_{\star}=70\,\,{M_{\odot}}$ and radius $R_{\rm star}=4\times 10^{10}\,\rm{cm}$ (Woosley & Heger, 2006). We adopt a (broken) power-law as the radial stellar mass density profile:
@@
where the density power-law indices $\alpha_{\rm p}$ of our simulations are listed in Tab. 1. The initial thermal pressure profile in the star is described by:
@@
where $P_{0}=0.1$ for models without and $P_{0}=0.25$ for models with constant density core. This prefactor is chosen such that the pressure on the entire grid is within the temperature validity range of the equation of state (§ 3.3). The initial thermal pressure is unlikely to impact the disk dynamics since the pressure in the disk will be dominated by energy dissipation due to turbulence associated with the accretion process. Consequently, the primary effect of the initial gas pressure is to provide additional support against gravity, potentially delaying the disk formation and the subsequent infall of stellar gas onto the accretion disk.
@@
where we adopt a stellar core radius, $R_{\rm core}=10^{8}\,\rm{cm}$. This configuration results in a nearly uniform vertical magnetic field at $r\lesssim R_{\rm core}$ that turns radial inside the star and closes near the stellar surface. The magnetic moment $\mu$ is chosen so that magnetization in the core is $\sigma={b^{2}}/{\rho}\lesssim 0.1$, corresponding to the gas to magnetic pressure ratio of $\beta=\max P_{\rm gas}/\max P_{\rm mag}\sim$ few.

### v2

# Magnetically-Driven Neutron-Rich Ejecta Unleashed: Global 3D Neutrino-General Relativistic Magnetohydrodynamic Simulations of Collapsars Probe the Conditions for $r$-process Nucleosynthesis

@@ Abstract
Collapsars – rapidly rotating stellar cores that form black holes – can power gamma-ray bursts (GRBs) and are proposed to be key contributors to the production of heavy elements in the Universe via the rapid neutron capture process ($r$-process). Previous neutrino-transport collapsar simulations have been unable to unbind neutron-rich material from the disk. However, these simulations have not included sufficiently strong magnetic fields and the black hole (BH), both of which are essential for launching mass outflows. We present $\nu$h-amr, a novel neutrino-transport general relativistic magnetohydrodynamic ($\nu$GRMHD) code, which we use to perform the first 3D global $\nu$GRMHD collapsar simulations. We find a self-consistent formation of a weakly magnetized dense accretion disk, which has sufficient time to neutronize. Eventually, substantial magnetic flux accumulates near the BH, becomes dynamically important, leads to a magnetically arrested disk (MAD), and unbinds some of the neutron-rich material. However, the strong flux also hinders accretion, lowers density, and increases neutrino cooling timescale, which prevents further disk neutronization. Typical collapsar progenitors with mass accretion rates, $\dot{M}\sim 0.1-1M_{\odot}/\rm{s}$, do not produce significant neutron-rich ($Y_{\text{e}}<0.25$) ejecta. However, we find that MADs at higher mass accretion rates, $\dot{M}\gtrsim\text{few}\,M_{\odot}/\rm{s}$ (e.g., for more centrally concentrated progenitors), can unbind $M_{\text{ej}}\lesssim{}M_{\odot}$ of neutron-rich ejecta. The outflows inflate a shocked cocoon that mixes with the infalling neutron-poor stellar gas and raises the final outflow $Y_{\text{e}}$; however, the final $r$-process yield may be determined earlier at the point of neutron capture freeze-out. Future work will explore under what conditions more typical collapsar engines become $r$-process factories.
@@ 1 Introduction
Generally, the $r$-process nucleosynthesis channel requires neutron-rich outflows. Accretion disks that form during the collapse of rapidly rotating massive stars exhibit similar density and temperature conditions as those found in postmerger disks. Moreover, the larger size and lifetime of collapsar disks render them more massive than postmerger disks, suggesting their potential as the major source of the Universe’s heavy $r$-process elements (Siegel et al., 2019). However, the stellar infall, which forms collapsar disks, likely obstructs the ejection of neutron-rich disk outflows. Describing this complex, nonlinear behavior necessitates numerical simulations to accurately model the formation of the collapsar disk, the synthesis of heavy elements, and the propagation of outflows through the collapsing star.
@@
The collapse of a massive stellar core initially forms a proto-NS. The $r$-process nucleosynthesis can happen in NS outflows, which are a combination of neutrino-driven (e.g., Thompson et al., 2004; Metzger et al., 2007) and magnetically driven (Kotake et al., 2004; Ardeljan et al., 2005; Moiseenko et al., 2006; Obergaulinger et al., 2006; Obergaulinger & Aloy, 2020, 2021; Mösta et al., 2014; Kuroda et al., 2020; Aloy & Obergaulinger, 2021). Neutrino-heated winds can facilitate favorable conditions for the $r$-process only for rapidly rotating, strongly magnetized proto-NSs, thanks to the shorter expansion timescales (Thompson et al., 2004; Metzger et al., 2007, 2008b; Vlasov et al., 2014, 2017; Desai et al., 2022; Prasanna et al., 2023). Studies of magnetorotational supernovae (MR-SNe) show that they require rapid rotation and strong magnetic fields to launch powerful magnetically driven outflows, or jets (Nishimura et al., 2006; Winteler et al., 2012; Nishimura et al., 2015; Tsujimoto & Nishimura, 2015; Halevi & Mösta, 2018; Mösta et al., 2018; Reichert et al., 2021, 2023). Strong outflows help to unbind neutron-rich matter and drive it at high velocities to avoid neutrino irradiation from the proto-NS, leading to robust $r$-process nucleosynthesis. Weaker initial magnetic field strengths lead to weaker outflows or delayed jet launching and therefore result in the $r$-process element production only up to the second peak (Nishimura et al., 2006; Winteler et al., 2012; Nishimura et al., 2015). Moreover, effects like current-driven kink instabilities and misalignment between the magnetic and rotation axes can also reduce the ability of MR-SNe to contribute to heavy $r$-process enrichment (Halevi & Mösta, 2018; Mösta et al., 2018).
@@
Alternatively, neutron-rich accretion disks can also form around BHs in collapsars (Woosley, 1993; MacFadyen & Woosley, 1999; MacFadyen et al., 2001; Fujimoto et al., 2004; Pruet et al., 2004; Kohri et al., 2005; Surman et al., 2006; Woosley & Bloom, 2006) or in common-envelope jetted explosions (Grichener & Soker, 2019). Growing computational power and advances in computational methods have enabled the numerical simulations of the $r$-process nucleosynthesis in this context. Earlier calculations showed neutron-rich material production and successful $r$-process operation in collapsar jets (Fujimoto et al., 2007; Ono et al., 2012; Nakamura et al., 2013). More recently, simulations modeled the magnetized evolution of an idealized isolated preset torus in hydrostatic equilibrium, i.e., without including the progenitor star (Siegel et al., 2019; Miller et al., 2020). These works show that collapsar disks may become neutron rich. However, it remains unclear whether the behavior of these idealized tori resembles that of disks self-consistently formed during the stellar collapse and whether the neutron-rich elements can overcome the ram pressure of the infalling stellar gas and escape from the collapsing star. More recently, simulations included the collapsing star for the first time, which hindered the ejection of the heavy $r$-process elements (Just et al., 2022; Fujibayashi et al., 2023; Dean & Fernández, 2024a, b). However, these simulations lacked important physics: they reduced the dimensionality to 2D and excluded magnetic effects. The absence of magnetic fields does not allow magnetized outflows to form via, e.g., the Blandford-Znajek (BZ; Blandford & Znajek, 1977) and Blandford-Payne–like (Blandford & Payne, 1982) processes, which may potentially carry away some of the neutron-rich elements.
@@
In this Letter, we present the first 3D neutrino-transport general-relativistic magnetohydrodynamic ($\nu$GRMHD) collapsar simulations. Our simulations use a two-moment (M1) neutrino-transport scheme. We find that including both the self-consistent disk formation in the collapsing stellar gas and subsequent magnetic launching of collimated disk outflows is crucial for the development of neutron-rich ejecta in collapsars. We begin by discussing the theoretical requirements for forming and ejecting neutron-rich material in §2. Using these arguments, we motivate our numerical setup in §3. We present our simulation results in §4 and discuss their implications in §5.
@@ 2 How to Make $r$-process Ejecta
Here, we outline the conditions under which collapsar disks form and generate neutron-rich material, and subsequently eject it. An accretion disk forms if the collapsing stellar gas possesses specific angular momentum, $l(r)$, exceeding that at the innermost stable circular orbit (ISCO), $r_{\rm ISCO}$:
@@
where $r$ is the spherical polar radius and $\,{M_{\rm BH}}$ is the BH mass, and for simplicity, we use a nonrelativistic approximation for the specific angular momentum at the ISCO. There are theoretical indications that collapsar progenitors feature rapid rotation, which results in early disk formation (e.g., Gottlieb et al., 2024).
A necessary, but not sufficient, condition for the $r$-process is that the disk must become neutron rich. The degree of neutron richness is inversely correlated with the electron fraction $\,{Y_{\rm e}}=n_{\rm p}/(n_{\rm p}+n_{\rm n})$, where $n_{\rm p}$ and $n_{\rm n}$ are the proton and neutron number densities, respectively. For weak interactions to drive the disk material neutron-rich ($\,{Y_{\rm e}}<0.5$), it must become dense and mildly degenerate through neutrino cooling (e.g., Beloborodov 2003). Specifically, when the viscous heating rate is counterbalanced by neutrino cooling, the disk transitions from a thick advective state to an efficiently neutrino-cooled regime (e.g., Di Matteo et al. 2002), in which it can become notably dense and neutron rich. Satisfying this “ignition” condition requires a minimum mass accretion rate at the ISCO (Chen & Beloborodov, 2007; Metzger et al., 2008a, b),
@@
where $\alpha_{\rm eff}$ is the effective viscosity and $K_{\rm ign}$ is a numerical prefactor that depends on the BH spin $a_{\rm BH}$: $K_{\rm ign}=0.07$ for $a_{\rm BH}=0$ and $K_{\rm ign}=0.02$ for $a_{\rm BH}=0.95$.
@@
and consequently becomes neutron rich. At the ignition radius, $R=R_{\rm ign}$, the neutrino cooling and viscous heating balance each other. In classical accretion disk theory (e.g., for a standard Shakura & Sunyaev 1973 disk), a single viscosity coefficient, $\alpha_{\rm eff}$, governs both the viscous heating and accretion. However, in the presence of strong magnetic fields, large-scale magnetic torques can render the accretion more efficient and decouple it from the heating (Blandford & Payne, 1982; Ferreira & Pelletier, 1995). In practice, for strongly magnetized disks, the accretion timescale is a factor of a few shorter than the viscous heating timescale (Manikantan et al., 2024). Therefore, we can define the ignition radius as the radius where the neutrino-cooling timescale, $t_{\rm cool}$, matches the accretion timescale, $t_{\rm acc}$:
@@
Here, the cooling timescale is (Metzger et al., 2008b)
@@
where $\dot{N}_{e^{-}p}$ is the rate of electron capture on protons and $\langle\cdot\rangle^{\theta,\varphi}_{\rho}$ is a density-weighted average over angles $\theta,\varphi$, which for quantity $X$ is
@@
Whereas disk winds may eject neutron-rich material in compact binary mergers (e.g., Fernández & Metzger, 2013; Siegel & Metzger, 2017; Radice et al., 2018), they might be too weak to unbind the neutron-rich material in the inner parts of collapsar disks (e.g., Dean & Fernández, 2024a). However, powerful relativistic BZ jets, associated with gamma-ray bursts (GRBs), can explode the star and eject the neutron-rich material out of the collapsing star. The formation of such jets becomes possible once enough poloidal (i.e., pointing in the $r$- and $\theta$-directions) magnetic flux accumulates on the BH to overcome the ram pressure of the infalling gas. Namely, when the Alfvén velocity becomes comparable to the free-fall velocity, the jet can emerge from the BH event horizon (Komissarov & Barkov, 2009; Gottlieb et al., 2023).
The jet-launching criterion is determined by the dimensionless magnetic flux on the BH, defined as
@@
where $\,{r_{\rm g}}\equiv GM_{\rm BH}/c^{2}$ is the BH gravitational radius. To infer the minimum value of $\phi$ for jet launching, we use estimates by Komissarov & Barkov (2009), where they found $\phi_{\mathrm{crit}}\sim 20$ in idealized simulations of jet launching for radial magnetic geometry (this corresponds to $\kappa_{\rm c}\sim 1.5$ in their work). In addition to powerful BH-powered outflows, magnetically arrested disks (MADs; Narayan et al., 2003; Tchekhovskoy et al., 2011) can also launch massive disk winds whose properties we study below. The accretion flow reaches the maximum magnetization when it goes to MAD, which occurs at $\phi\approx 50$ (Tchekhovskoy et al., 2011).
We conclude that collapsars can eject $r$-process elements if (1) they form an accretion disk with high mass accretion rates and efficient neutrino cooling, and (2) once the $r$-process elements are synthesized in the disk, dynamically important magnetic fields around the BH launch magnetized jets and winds that manage to unbind the neutron-rich material.
@@ 3.1 Initial conditions
Table 1: Collapsar model parameters. The leftmost column: model name, which follows the [cp]B[0ws][A#] convention, where c or p indicates stellar density profile (core or power law, respectively); 0, w, or s indicates the magnetic field strength (zero, weak, and strong field); and A# indicates BH spin (omitted for the fiducial value, $a_{\rm BH}=0.8$). The rest of the columns, from left to right: initial density slope, $\alpha_{\rm p}$; grid resolution, $N_{r}\times N_{\theta}\times N_{\phi}$; simulation duration, $t_{\rm f}$; maximum initial core magnetic field strength, $B_{\rm core}$; MAD onset time, $t_{\rm MAD}$; jet (outflow) launch time, $t_{\rm jet}$; $\dot{M}$ at these two times; jet power at the larger of these two times, $L_{\rm j}$; and total mass of neutron-rich ($\,{Y_{\rm e}}<0.25$) material at $t_{\rm MAD}$, $M_{\rm nr}$. Parameter values common to all models are BH mass, $M_{\rm BH}=4\,{M_{\odot}}$; stellar envelope total mass $M_{\rm star}=70\,{M_{\odot}}$; stellar radius, $R_{\rm star}=4\times 10^{10}$ cm; and magnetic core radius, $R_{\rm core}=10^{8}$ cm (such that $\mu=B_{\rm core}R_{\rm core}^{2}/2$ in Eq. 12).
@@
To determine the electron fraction evolution, we implement a novel two-moment (M1) neutrino-transport scheme alongside the Helmholtz equation of state (EOS; see §3.4) into the GPU-accelerated code h-amr (Liska et al., 2022). Using the resulting $\nu$h-amr code, we conduct a suite of 3D $\nu$GRMHD simulations of collapsing stars, as outlined in Table 1. In the following, we adopt the units $G=c=1$ and use Heaviside-Lorentz units for magnetic fields, thereby absorbing the factor of $1/\sqrt{4\pi}$ into the definition of magnetic field strength, $B$.
We set up our collapsar simulations following Gottlieb et al. (2022a, b). We set the newly formed BH mass to $\,{M_{\rm BH}}=4\,\,{M_{\odot}}$ and its dimensionless spin magnitude to $\,{a_{\rm BH}}=0.8$. Our simulation incorporates a static metric such that the BH mass and spin do not evolve over the course of the simulation. The BH is embedded in a stripped-envelope (e.g., Wolf–Rayet-like) star of mass [^1] As we will show, at the relevant simulation times (MAD onset timescale), the accreted mass is $\lesssim 10\,\,{M_{\odot}}$. This implies that a slightly more extended core with a steeper density profile outside will produce the same accretion rates with $M_{\star}\approx 20\,\,{M_{\odot}}$. $M_{\star}=70\,\,{M_{\odot}}$ and radius $R_{\rm star}=4\times 10^{10}\,\rm{cm}$ (Woosley & Heger, 2006). We adopt a (broken) power law as the radial stellar mass density profile:
@@
where the density power-law indices $\alpha_{\rm p}$ of our simulations are listed in Table 1. We simulate two progenitor models: a typical collapsar with $\alpha_{\rm p}=1.5$, representing the density profile established in the aftermath of BH formation (Halevi et al., 2023); and a collapsar with an extreme, more centrally concentrated distribution with a constant density core of radius $10^{8},\rm{cm}$, followed by a density slope of $\alpha_{\rm p}=2.5$, to probe neutron-rich outflow ejection in the regime of high mass accretion rates.
The initial thermal pressure profile in the star is described by
@@
where $P_{0}=0.1$ for models without and $P_{0}=0.25$ for models with constant density core. This prefactor is chosen such that the pressure on the entire grid is within the temperature validity range of the EOS (§ 3.3). The initial thermal pressure is unlikely to impact the disk dynamics since the pressure in the disk will be dominated by energy dissipation due to turbulence associated with the accretion process. Consequently, the primary effect of the initial gas pressure is to provide additional support against gravity, potentially delaying the disk formation and the subsequent infall of stellar gas onto the accretion disk.
@@
where we adopt a stellar core radius, $R_{\rm core}=10^{8}\,\rm{cm}$. This configuration results in a nearly uniform vertical magnetic field at $r\lesssim R_{\rm core}$ that turns radial inside the star and closes near the stellar surface. The magnetic moment $\mu$ is chosen so that magnetization in the core is $\sigma={b^{2}}/{\rho}\lesssim 0.1$, corresponding to the gas-to-magnetic-pressure ratio of $\beta=\max P_{\rm gas}/\max P_{\rm mag}\sim$ few.